# Breaking Imitation Bottlenecks: Reinforced Diffusion Powers Diverse Trajectory Generation

### v1

@@ Abstract
Most end-to-end autonomous driving methods rely on imitation learning from single expert demonstrations, often leading to conservative and homogeneous behaviors that limit generalization in complex real-world scenarios. In this work, we propose DIVER, an end-to-end driving framework that integrates reinforcement learning with diffusion-based generation to produce diverse and feasible trajectories. At the core of DIVER lies a reinforced diffusion-based generation mechanism. First, the model conditions on map elements and surrounding agents to generate multiple reference trajectories from a single ground-truth trajectory, alleviating the limitations of imitation learning that arise from relying solely on single expert demonstrations. Second, reinforcement learning is employed to guide the diffusion process, where reward-based supervision enforces safety and diversity constraints on the generated trajectories, thereby enhancing their practicality and generalization capability. Furthermore, to address the limitations of L2-based open-loop metrics in capturing trajectory diversity, we propose a novel Diversity metric to evaluate the diversity of multi-mode predictions. Extensive experiments on the closed-loop NAVSIM and Bench2Drive benchmarks, as well as the open-loop nuScenes dataset, demonstrate that DIVER significantly improves trajectory diversity, effectively addressing the mode collapse problem inherent in imitation learning.
@@ 4.1 Overall Architecture
As shown in Figure 3, we propose DIVER, a novel E2E-AD framework that integrates diffusion models and reinforcement learning to enable diverse, and safe multi-mode trajectories. DIVER consists of two main components: perception module and motion planner, which takes raw perception features as input and outputs a distribution over future trajectories. At the training paradigm level, DIVER adopts a hybrid learning scheme that combines IL with RL. The core innovations of DIVER lie in two aspects. First, the Policy-Aware Diffusion Generator (PADG) generates diverse trajectory candidates by leveraging a diffusion model conditioned not only on a single expert trajectory but also on multiple Reference GTs, capturing a range of plausible human-like driving behaviors such as lane changes, yielding, and overtaking. Second, we leverage reinforcement learning to guide the diffusion-based trajectory generation process, addressing the inherent limitations of IL-based supervision and aligning the output with safety and diversity objectives. Overall, DIVER combines the generative capability of diffusion models with the optimization strength of reinforcement learning, forming a flexible and scalable framework for E2E-AD. It enables the system to explore a broader solution space in complex scenarios while ensuring the physical feasibility and behavioral safet of the generated trajectories.
@@
To overcome the limitations of conservative and mode-collapsed behaviors in IL-based E2E-AD, we propose the Policy-Aware Diffusion Generator (PADG). PADG is a core component of our DIVER that enables diverse, and feasible trajectory generation through a conditional denoising diffusion process. Specifically, PADG is built upon a conditional diffusion framework, which reconstructs multi-mode future trajectories distributions from random noise through iterative denoising, guided at each step by rich scene semantics as conditional information. It consists of a dual-branch architecture: one branch learns to reconstruct the distribution of the predicted trajectory, while the other models the expert trajectory to serve as guidance.
@@
4.2.2 Conditional Diffusion Feature Interaction.
@@
In the first stage, we apply a trajectory-aware image feature pooling module, referred to as TrajPooler (Algorithm 1). Given a noisy trajectory $\hat{\tau}^{(m)}$, we first recover its absolute coordinates via cumulative sum. We then sparsely sample 3D points along the trajectory and project them onto the image plane using the known camera intrinsics $\mathbf{K}$ and extrinsics $\mathbf{T}$. These projected 2D locations are used to extract multi-scale image features from the BEV encoder’s output. To adaptively aggregate these image features, we learn attention weights $\alpha^{(m)}$ conditioned on the instance-level query $\mathcal{F}{\text{inst}}^{(m)}$. The fused trajectory feature $F{\text{traj}}^{(m)}$ is obtained via a weighted sum followed by an MLP and residual fusion with the original instance embedding.
@@
Instance feature $\mathcal{F}_{\text{inst}}^{(m)}\in\mathbb{R}^{d}$,
@@
6Learned attention weights: $\alpha^{(m)}\leftarrow\text{AttentionWeights}(\mathcal{F}_{\text{inst}}^{(m)},\mathbf{K},\mathbf{T})$
Fused trajectory feature: $F_{\text{traj}}^{(m)}\leftarrow\text{MLP}(\sum\alpha^{(m)}\cdot F_{2D}^{(m)})+\mathcal{F}_{\text{inst}}^{(m)}$
@@ 4.3.1 Overview of Group Relative Policy Optimization
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{\text{old}}}(a|s)}\cdot\left(\hat{A}(s,a)-\mathbb{E}_{g}[\hat{A}(s,a)]\right)\right],$ (16)
where $\hat{A}(s,a)$ is the estimated advantage, and $\mathbb{E}_{g}[\cdot]$ denotes the expectation over a group of trajectories or modes. This relative formulation enhances training stability while promoting policy diversity.
@@ 4.3.2 GRPO for Trajectory Planning
We define a composite reward function $r(\hat{\tau})$ that evaluates each trajectory in terms of diversity and safety. The detailed formulations of the Diversity Reward $r_{\text{div}}$ and Safety Reward $r_{\text{safe}}$ are presented below.
@@ 5.2 Evaluation Metrics
[…]
@@ 5.3 Implementation Details
[…]
@@
${\operatorname{DIVER}}^{*}$ (Ours) MT IL&RL - 1.11 0.38 68.90 36.75 72.34 22.34 35.08 25.09 41.09 50.00 59.21 42.09
@@
[…]
@@
[…]
@@
$\mathcal{L}_{\text{match}}$+$\mathcal{L}_{\text{RL(PPO)}}$ 0.09 0.19 0.32 0.20
$\mathcal{L}_{\text{match}}$+$\mathcal{L}_{\text{RL(GRPO)}}$ 0.10 0.19 0.34 0.21
@@
$\mathcal{L}_{\text{match}}$+$\mathcal{L}_{\text{RL(PPO)}}$ 98.1 95.4 94.1 100 82.3 87.9
$\mathcal{L}_{\text{match}}$+$\mathcal{L}_{\text{RL(GRPO)}}$ 98.5 96.5 94.9 100 82.6 88.3
@@
To demonstrate the generalization capability of DIVER, we conduct experiments against several strong E2E-AD baselines, including SparseDrive  [3], VAD  [2], and TransFuser [5]. On the Bench2Drive  [50], nuScenes  [51], and Adv-nuSc  [59] datasets, we compare against SparseDrive and VAD. For nuScenes and Adv-nuSc, we adopt ResNet-50  [68] as the image backbone with an input resolution of 256$\times$704. The detection module operates within a circular range of 55 meters, and the online mapping covers a 60$\times$30 meter area (longitudinal×$\times$×lateral). The motion module generates 6 trajectory candidates (modes). On Bench2Drive, we use ResNet-50 with 6 decoder layers and an input resolution of 640$\times$352. We also define fixed numbers of hybrid task queries: 900 agent queries, 100 map queries, and 480 planning queries. On the NAVSIM  [55] dataset, we adopt TransFuser  [5] as the baseline and follow the standard Navtrain split for training. For fair comparison, we use the same perception modules and ResNet-34 backbone as in TransFuser. All experiments are conducted on a server with 8 NVIDIA RTX 4090 (24GB) GPUs.
@@ 5.6.1 Ablation Study on PADG
[…]
@@ 5.6.2 Ablation Study on Loss Function Design
[…]

### v2

@@ Abstract
Existing end-to-end autonomous driving (E2E-AD) methods predominantly rely on single expert demonstrations through imitation learning, often leading to conservative and homogeneous driving behaviors that struggle to generalize to complex real-world scenarios. In this work, we propose DIVER, a novel E2E-AD framework that combines diffusion-based multi-mode trajectory generation with reinforcement learning to produce diverse, safe, and goal-directed trajectories. First, the model conditions on map elements and surrounding agents to generate multiple reference trajectories from each ground-truth reference trajectory that overcome the inherent limitations of single-mode imitation. Second, we treat the diffusion process as a stochastic policy and employ Group Relative Policy Optimization (GPRO) objectives to guide the diffusion process. By optimizing trajectory-level rewards for both diversity and safety, GRPO directly mitigates mode collapse and enhances collision avoidance, encouraging exploration beyond expert demonstrations and ensuring physically plausible plans. Furthermore, to address the limitations of L2-based open-loop metrics in capturing trajectory diversity, we propose a novel trajectory diversity metric to evaluate the diversity of multi-mode predictions. Extensive experiments on the closed-loop NAVSIM and Bench2Drive benchmarks, as well as the open-loop nuScenes dataset, demonstrate that DIVER significantly improves trajectory diversity, effectively addressing the mode collapse problem inherent in imitation learning.
@@ 4.1 Overall Architecture
As shown in Figure 3, we propose DIVER, a novel E2E-AD framework that integrates diffusion models and reinforcement learning to enable diverse, and safe multi-mode trajectories. DIVER consists of two main components: perception module and motion planner, which takes raw perception features as input and outputs a distribution over future trajectories. At the training paradigm level, DIVER adopts a hybrid learning scheme that combines IL with RL. The core innovations of DIVER lie in two aspects. First, the Policy-Aware Diffusion Generator (PADG) generates diverse trajectory candidates by leveraging a diffusion model conditioned not only on a single expert trajectory but also on multiple Reference GTs, capturing a range of plausible human-like driving behaviors such as lane changes, yielding, and overtaking. Second, we leverage reinforcement learning to guide the diffusion-based trajectory generation process, addressing the inherent limitations of IL-based supervision and aligning the output with safety and diversity objectives. Overall, DIVER combines the generative capability of diffusion models with the optimization strength of reinforcement learning, forming a flexible and scalable framework for E2E-AD. It enables the system to explore a broader solution space in complex scenarios while ensuring the physical feasibility and behavioral safety of the generated trajectories.
@@
To overcome the limitations of conservative and mode-collapsed behaviors in IL-based E2E-AD, we propose the Policy-Aware Diffusion Generator (PADG). PADG is a core component of our DIVER that enables diverse, and feasible trajectory generation through a conditional denoising diffusion process. Specifically, PADG is built upon a conditional diffusion framework, which reconstructs multi-mode future trajectory distributions from random noise through iterative denoising, guided at each step by rich scene semantics as conditional information. It consists of a dual-branch architecture: one branch learns to reconstruct the distribution of the predicted trajectory, while the other models the expert trajectory to serve as guidance.
@@
4.2.2 Conditional Diffusion Decoder.
@@
In the first stage, we apply a trajectory-aware image feature pooling module, referred to as TrajPooler (Algorithm 1). Given a noisy trajectory $\hat{\tau}^{(m)}$, we first recover its absolute coordinates via cumulative sum. We then sparsely sample 3D points along the trajectory and project them onto the image plane using the known camera intrinsics $\mathbf{K}$ and extrinsics $\mathbf{T}$. These projected 2D locations are used to extract multi-scale image features from the BEV encoder’s output. To adaptively aggregate these image features, we learn attention weights $\alpha^{(m)}$ conditioned on the initial trajectory feature $F{\text{traj}}^{(m)}$. The fused trajectory feature $F{\text{traj}}^{(m)}$ is obtained via a weighted sum followed by an MLP and residual fusion with the original instance embedding.
@@
Trajectory feature $F_{\text{traj}}^{(m)}\in\mathbb{R}^{d}$,
@@
6Learned attention weights: $\alpha^{(m)}\leftarrow\text{AttentionWeights}(F_{\text{traj}}^{(m)},\mathbf{K},\mathbf{T})$
Fused trajectory feature: $F_{\text{traj}}^{(m)}\leftarrow\text{MLP}(\sum\alpha^{(m)}\cdot F_{2D}^{(m)})+F_{\text{traj}}^{(m)}$
@@ 4.3.1 Overview of Group Relative Policy Optimization
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{q,\left\{o_{i}\right\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{KL}\left(\pi_{\theta}\|\pi_{\text{ref }}\right)\right],$ (16)
$\mathcal{L}_{i}=\min\left(w_{i}A_{i},\operatorname{clip}\left(w_{i},1-\epsilon,1+\epsilon\right)A_{i}\right),$ (17)
where, $w_{i}=\frac{\pi_{\theta}\left(o_{i}\mid q\right)}{\pi_{\theta_{old}}\left(o_{i}\mid q\right)}$, $\epsilon$ and $\beta$ are hyper-parameters, and the advantage $A_{i}$ is computed using the normalized reward within the group.
@@ 4.3.2 GRPO for Trajectory Planning
[…]
@@ 5.2 Evaluation Metrics
$\textit{Div.}^{(t)}=\min\left(1,\frac{D_{\text{raw}}^{(t)}}{\epsilon+\frac{1}{M}\sum_{m=1}^{M}\left\|\mathbf{p}_{t}^{(m)}\right\|_{2}}\right),$ (24)
@@ 5.3 Implementation Details
TABLE I: $\operatorname{\textbf{Open-Loop}}$, $\operatorname{\textbf{Closed-Loop}}$ results and $\operatorname{Multi-Ability}$ results on $\operatorname{Bench2Drive}$ (V0.0.3) under base training set. ‘mmt’ refers multi-mode trajectories variant of $\operatorname{VAD}$ and denotes the re-implementation. * denotes expert feature distillation. ‘DS’ denotes Driving Score. ‘SR’ denotes Success Rate. ‘Effi’ denotes Efficiency. ‘Comf’ denotes Comfortness. ‘Merg.’ denotes Merging. ‘Overta.’ denotes Overtaking. ‘Emerge.’ denotes Emergency Brake. It is worth noting that the Diversity Metric ($\operatorname{\textit{Div.}^{(t)}}\uparrow$) is applicable only to multi-mode E2E-AD methods.
@@
$\operatorname{DIVER}$} (Ours) MT IL&RL - 1.11 0.38 68.90 36.75 72.34 22.34 35.08 25.09 41.09 50.00 59.21 42.09
@@
TABLE II: Comparison on planning-oriented $\operatorname{NAVSIM}$ navtest split with $\operatorname{\textbf{Closed-Loop}}$ metrics. ‘mmt’ refers multi-mode trajectories variant of $\operatorname{TransFuser}$ and denotes the re-implementation.
@@
TABLE III: Planning results on the $\operatorname{\textbf{Open-Loop}}$ $\operatorname{NuScenes}$ [51] validation dataset. It is worth noting that the Diversity Metric ($\operatorname{\textit{Div.}^{(t)}}\uparrow$) is applicable only to multi-mode E2E-AD methods like MomAD [58], DiffusionDrive [4], and SparseDrive [3].
@@
$\mathcal{L}_{\text{match}}$$\mathcal{L}_{\text{RL(PPO)}}$ 0.09 0.19 0.32 0.20
$\mathcal{L}_{\text{match}}$$\mathcal{L}_{\text{RL(GRPO)}}$ 0.10 0.19 0.34 0.21
@@
$\mathcal{L}_{\text{match}}$$\mathcal{L}_{\text{RL(PPO)}}$ 98.1 95.4 94.1 100 82.3 87.9
$\mathcal{L}_{\text{match}}$$\mathcal{L}_{\text{RL(GRPO)}}$ 98.5 96.5 94.9 100 82.6 88.3
@@
[…]
@@ 5.6.1 Ablation Study on PADG
Effect of Condition and Reference GTs. As shown in Tables VIII, IX, and X, our ablation studies on Bench2Drive[50], NAVSIM[55], and NuScenes[51] validate the effectiveness of the PADG in enhancing both diversity and safety in trajectory generation. Condition refers to the incorporation of map and agent information into the diffusion process, enabling more context-aware and structurally grounded predictions. This leads to improved multi-modeity (14.3% increase in $\operatorname{\textit{Div.}^{(t)}}$ on Bench2Drive) and enhanced planning safety and comfort (NC +0.1%, PDMS +1.2 on NAVSIM). Reference GTs, on the other hand, introduce multiple high-quality trajectory anchors to guide the diffusion process beyond a single expert demonstration. This mitigates mode collapse and encourages diverse yet feasible trajectory generation. The integration of Reference GTs further boosts closed-loop performance (DS +1.78, SR +1.57% on Bench2Drive) and improves navigation robustness (TTC +0.8 on NAVSIM). When both components are combined in DIVER, we observe consistent state-of-the-art results across all benchmarks. These results demonstrate that context conditioning and reference guidance play complementary roles in generating diverse, safe, and realistic driving behaviors.
@@ 5.6.2 Ablation Study on Loss Function Design
As shown in Tables XI and XII, loss function design has a significant impact on model performance. On the NuScenes dataset, the standard L1 loss fails to exploit the potential of diffusion and reference GTs, yielding limited diversity (Avg. Div.↑ = 0.14). Even when combined with RL (PPO), improvements are marginal. In contrast, replacing L1 with the proposed matching loss $\mathcal{L}_{\text{match}}$ enables the model to benefit from trajectory diversity, yielding higher $\operatorname{\textit{Div.}^{(t)}}$ (up to 0.21). Further integrating RL with diversity and safety rewards enhances performance, especially using GRPO, which outperforms PPO in both diversity and planning quality. On NAVSIM, $\mathcal{L}_{\text{match}}+\mathcal{L}_{\text{RL(GRPO)}}$ achieves the best PDMS (88.3), confirming that this design effectively balances realism, diversity, and safety.